Explain the essential features of a refracting telescope. Define and deduce an expression for the magnifying power of this instrument.


1
Expert's answer
2021-03-16T11:39:13-0400

Answer

Magnifying power of telescope is the ratio of the angle "\alpha" subtended at the eye by the final imege to angle "\beta" which the object subtends at the lenses.


"m=\frac{\beta}{\alpha}=\frac{h}{f_e}\frac{f_o}{h}\\m=\frac{f_o}{f_e}"